# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v10

@@ Abstract
I analyzed the risk-coping strategies among factory worker households in early 20th-century Tokyo. I digitized and analyzed a unique daily longitudinal household budget survey to determine how consumption was impacted by idiosyncratic shocks. I found that while the households were so vulnerable that the shocks impacted their consumption levels, the estimated income elasticity for food consumption was relatively low in the short run. The event-study analysis using adverse health shock confirms the robustness of the results. The result of mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption. Despite the potential loss of profit, retailers in a competitive situation allowed consumers to trade on credit. This shows the roles of informal credit institutions in mitigating vulnerability among urban worker households.
@@ 1 Introduction
Economic theory illustrates the mechanism behind the smoothing of consumption in complete markets (Arrow and Debreu 1954). Additionally, a series of empirical studies have tested the smoothness of consumption among households in various economies under systematic empirical specification. Evidence shows that markets in contemporary developing countries are often imperfect.111 The early representative literature includes, for example, Townsend (1994), Townsend (1995), and Ravallion and Chaudhuri (1997). Morduch (1999), Townsend (1995), and Dercon (2004) provide reviews of the earlier studies. Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions. The households are vulnerable to risk, and idiosyncratic shocks impact the levels of their consumption because shocks are not fully insured in the markets. Therefore, consumption smoothing is an essential mechanism for mitigating poverty, especially in developing economies. [^2] Smoothness of consumption is also an important factor in human capital formation. For example, Foster (1995) investigates the shock of the 1988 floods in Bangladesh and shows that children’s growth patterns are influenced by the inefficiency of the credit markets. Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall events in early life increased the survival probabilities of girls in rural India. Gertler and Gruber (2002) also shows that consumption smoothing against health shocks is imperfect in Indonesia.
Credit institutions in local society can mitigate idiosyncratic income shocks, especially when the availability of formal financial institutions is limited. Udry (1994) showed that state-contingent loans among relatives and neighbors could partially insure against idiosyncratic shocks on rural households in northern Nigeria. A detailed field survey revealed that asymmetric information was no longer important in those private credit contracts (Udry 1990). Fafchamps and Lund (2003) also presented evidence that the households in the rural Philippines used informal networks with relatives and friends as a risk-mitigation strategy for idiosyncratic shocks. Related studies consider risk-sharing among households with similar cultural and institutional backgrounds. Grimard (1997) showed that insurance among households in the same ethnic group could partially mitigate idiosyncratic shocks in the multiethnic economy of Côte d’Ivoire. Munshi and Rosenzweig (2016) presented the importance of rural insurance networks among workers in the same caste in the context of spatial wage inequality in India. The evidence suggests that the availability of informal credit institutions relates to the stability of consumption in rural economies.
@@
This study assesses risk-coping behavior in informal credit transactions in the urban community in early twentieth-century Tokyo. I digitized and analyzed a unique daily longitudinal budget survey on factory worker households conducted in a representative manufacturing area. I found that the full insurance hypothesis is strongly rejected for overall consumption expenditure, meaning that the factory worker households in the historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks. However, expenses on indispensable consumption categories were partially smoothed in the short run. Credit purchases in local retailers played an important role in mitigating adverse idiosyncratic shocks, whereas the other risk-mitigation strategies were impractical. Credit purchases did not require an explicit contract and thus often led to a loss in sales for the retailers. However, to remain competitive, retailers allowed credit transactions for their customers to remain regular customers. This shows that informal credit institutions in urban local communities functioned as informal insurance networks in rural economies.
The first contribution of this study is to offer a case study of a metropolis in a developing economy. Urbanization is now a worldwide phenomenon. Moreover, the proportions of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020). This requires a deeper understanding of consumption behavior among urban working households in the economic development process. However, risk-mitigation strategies among urban worker households have been understudied. [^3] Extant studies have focused particularly on the rural economies (e.g., Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002). Recent field experimental studies also focus on village economies (Comola and Prina 2023; Macours et al. 2022). Although several studies considered consumption smoothing among urban households in postwar Bangkok and prewar Osaka, the risk-mitigation mechanisms, particularly via credit institutions, are not fully evaluated. [^4] See Townsend (1995) and Ogasawara (2024), respectively. Osaka was middle scale city of approximately half the size of Tokyo City circa 1920 (Online Appendix Table A.1). (Skoufias and Quisumbing 2005; Aguila et al. 2017) also provided case studies on middle-sized contemporary economies. However, they did not consider the roles of local credit institutions. This study considers a set of homogeneous households in a representative manufacturing area in prewar Tokyo, which provides a valuable empirical setting for testing risk-sharing strategies in the historical metropolis. [^5] Focusing on a sample that shares similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). My findings show that informal credit institutions in the local urban community could mitigate vulnerability among households similar to rural developing economies (Udry 1994; Fafchamps and Lund 2003).
@@
The second contribution is to add analytical perspectives to economic history studies. First, studies of how working-class households overcame their economic hardships during the late 19th to the early 20th century have focused primarily on Britain. The weight of evidence supports the roles of historical financial institutions and labor supply adjustments in compensating for income loss in working-class households. [^6] Horrell and Oxley (2000) highlight the important role of labor supply adjustments among family members to manage the income losses of the head in working-class households in late 19th-century Britain. Scott and Walker (2012) also found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use savings institutions. Kiesling (1996) and Boyer (1997) focused on the risk-mitigation strategies among the cotton workers households in Lancashire during the recession caused by the Civil War. However, there is little evidence of how those risk-coping strategies smoothed their consumption. Similarly, previous studies have used cross-sectional variations in income to estimate the elasticities of the temporary income sources. [^7] For example, Horrell and Oxley (2000) and James and Suto (2011) consider the deviations of actual earnings from the predicted earnings as the shocks. Scott and Walker (2012) employed the cross-sectional variations in the pooled cross-sectional sample to analyze the smoothness of expenditure (deviation from mean expenditure). An exceptional study is Saaritsa (2011), which shows the estimated elasticity of temporary income from a fixed effect one-way error component model using a panel household budget survey in 1928 Helsinki. The estimated elasticity is, therefore, potentially disturbed by household-specific unobservable confounding factors such as preference and learning ability. As risk-mitigation is a concept in the dynamic behavior of a specific household, the estimator derived from within variations of the units is more suitable for evaluating the roles of risk-mitigation strategies. This study implements a stylized econometric method for the panel dataset to offer comparable estimates for future studies. Second, this study adds micro-level evidence on consumer strategies during industrialization. Francks (2009) provides a comprehensive view of the long-run consumption history in Japan from the Edo era to the twentieth century. [^8] Several business history studies analyze the relationship between consumption and production. A representative work is Gordon (2013), which investigated the expansion of the sewing machine from many perspectives, including production, distribution, and gender social norms in Japan. However, these studies have not focused on the households’ risk-mitigation strategies. Some studies investigated the relationships between household chores and consumption strategies in interwar households (Tanimoto 2011, 2012). This study employs an alternative analytical view by focusing on short-run consumption strategies. An examination of how, and the extent to which, the local society could ensure household consumption provides valuable insights into consumption history literature. Third, several financial historical studies provide supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005). This study is the first to document the function of informal credit transactions by local retailers in the context of consumption smoothing. Munshi and Rosenzweig (2016) linked spatial wage inequality and men’s mobility in India with the availability of risk-mitigation networks among workers in the same caste. From this perspective, lower barriers of local credit institutions may motivate rural workers to urbanize, which may partially explain why Tokyo grew so fast circa WWI. This hypothesis requires testing in future research.
@@ 2.2 Tsukishima: A Representative Manufacturing Area
The population in Tokyo City has been growing since the end of the 19th century, and industrialization due to WWI spurred its growth. Tokyo City’s population share increased from $2.5$% in 1889 to approximately $4$% by 1920.141414In 1920, there were more than $10,000$ municipalities in Japan, suggesting how large Tokyo was. There was a difference of about 1 million between Tokyo and Osaka, the second-largest city by population. Online Appendix Figure A.1a summarizes the trends in the number of people in the top five cities between 1889 and 1920. Throughout the 1910s, the percentage of industrial and commercial workers increased in Tokyo City, building the city as a gigantic commercial and industrial place (Tanimoto 2002, p. 8). Column 1 of Panel A in Table 1 shows that nearly half of the male workers in Tokyo worked in the industrial sector as of 1920.
@@
Panel B in Table 1 summarizes the population and occupational statistics from the census to examine the characteristics of working-class households in Tsukishima. First, the average household size in Tsukishima is $4.3$—similar to but slightly smaller than the values in Tokyo City ($4.6$) and Kyobashi Ward ($4.7$). The percentage of the married male population is $40$%, which is similar to the range of the figures for Tokyo ($38$%) and Kyobashi ($36$%). The average age of men in Tsukishima is $25.1$, slightly lower than the figures for Tokyo and Kyobashi. While the sex ratio by age group in Tsukishima shows a similar trend to that of Tokyo and Kyobashi, the ratios for the 1414141419191919 and 2020202029292929 age bins are higher than those. Overall, these statistics reflect the fact that Tsukishima was a typical industrial area and many young male workers were employed there. The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory worker households. The percentage of married men in Tsukishima suggests that a certain proportion of the households in their early 20s might have been couples before they started raising children.
@@ 3.2 Sample Characteristics
The estimation strategy used for testing consumption-smoothing and risk-coping strategies requires panel structures for aggregate income and expenditure (Section 4). I use the semi-monthly series for baseline analysis, given that the head’s income was paid twice monthly. [^31] Section 3.3 provides details on the definitions of aggregate income and expenditure. Among the $40$ THBS households, there are five cross-sectional units in our semi-monthly data. I excluded these five cross-sectional units, leaving $35$ households. Next, I removed a household with incomplete income information. Finally, I excluded a household for which no information on payments to the credit purchases was recorded. Consequently, my THBS sample for the semi-monthly series included $33$ households with panel structures from January 12 to November 11, 1919.323232I also used the adjusted monthly panel dataset for sub-analysis in Section 4. Following the same trimming steps, my adjusted monthly panel dataset included $26$ households for the same period. Note that several units have only three semi-month cells, for example, less than two adjusted months. These units have a panel structure in the semi-monthly panel dataset, but not in the adjusted monthly panel. Online Appendix B.3 summarizes the finer details of the original data structure and trimming.
@@ 3.3 Aggregation
To manage the mismatch, I aggregated the daily observations between the 12th and 26th for the first half and those between the 27th and 11th of the next month for the second half. [^41] This means that the number of days included in a semi-month ranges from 13 (i.e., in February) to 16 (e.g., January). Note that this heterogeneity in the number of days over different semi-months is a cross-sectional constant. Thus, these differences are entirely captured using the semi-month fixed effects in the regression. This range is based on the fact that paydays in the second half of a month measured in the THBS dataset ranged from 27th to 31st. [^42] Note that I do not use the definition of semi-month starting from the 2nd (to the 16th) in each month because this method of stratification places the payday at the end of each semi-month cell. Online Appendix B.7 provides the evidence that while the definition using 2nd in each month as the threshold provides an improved series in income, the expenditure cannot chase the income correctly because the expenses measured under this definition are based on the income in the previous semi-month cell. My preferred definition was using the 12th of each month as this threshold offers a clean series of income and expenditure. Figure 4c shows the time-series plots of the income and expenditure from this adjusted semi-monthly panel dataset, whereas Figure 4e illustrates the unadjusted (calender) semi-monthly series using the 15th of each month as the threshold of the semi-month. Clearly, Figure 4e provides very rough measures of income and expenditure due to the measurement errors from the maladjustment of paydays. The most significant improvement in Figure 4c is that the expenditure can chase the income in each month, reflecting the households’ consumption strategies based on their income measured in the approximate exact timing. Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which is unavailable under the unadjusted series (Figure 4f). In Section 4.1, I will demonstrate how these measurement errors in the unadjusted dataset attenuate the estimates.

### v11

@@ Abstract
I analyzed the risk-coping strategies among factory worker households in early 20th-century Tokyo. I digitized and analyzed a unique daily longitudinal household budget survey to determine how consumption was impacted by idiosyncratic shocks. I found that while the households were so vulnerable that the shocks impacted their consumption levels, the estimated income elasticity for food consumption was relatively low in the short run. The analysis using adverse health shock confirms the robustness of the results. The result of mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption. Despite the potential loss of profit, retailers in a competitive situation allowed consumers to trade on credit. This shows the roles of informal credit institutions in mitigating vulnerability among urban worker households.
@@ 1 Introduction
Economic studies that analyze the relationship between historical economic development and financial institutions have focused on the alleviation of constraints on corporate financing through official financial systems (Acemoglu et al. 2005; Levine 2005). However, systematic research on the association between household financial institutions and consumption smoothing during industrialization remains limited. How did working-class households respond to idiosyncratic shocks and smooth their consumption before public insurance was established? Development economics research has confirmed that consumption smoothing is an essential mechanism for mitigating poverty and explaining human capital formation. [^1] Evidence shows that markets in contemporary developing countries are often imperfect. Thus, the households are vulnerable to risk, and idiosyncratic shocks impact the levels of their consumption. The early representative literature includes, for example, Townsend (1994), Townsend (1995), and Ravallion and Chaudhuri (1997). Morduch (1999), Townsend (1995), and Dercon (2004) provide reviews of the earlier studies. Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions. Since shocks are not fully insured in the markets, smoothness of consumption is also an important factor in human capital formation. Foster (1995) investigates the shock of the 1988 floods in Bangladesh and shows that children’s growth patterns are influenced by the inefficiency of the credit markets. Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall events in early life increased the survival probabilities of girls in rural India. Gertler and Gruber (2002) also shows that consumption smoothing against health shocks is imperfect in Indonesia. Surprisingly, however, economic history studies have not yet provided an answer to this important question. We still know little about how historical households maintained their livelihoods during industrialization.
@@
This study assesses risk-coping behavior among the working-class households in early twentieth-century Tokyo. I digitized and analyzed a unique daily longitudinal budget survey on factory worker households conducted in a representative manufacturing area. The full insurance hypothesis is strongly rejected for overall consumption expenditure, indicating that factory worker households in the historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks. However, expenses on indispensable consumption categories were partially smoothed in the short run. Credit purchases in local retailers played an important role in mitigating adverse idiosyncratic shocks, whereas the other risk-mitigation strategies were impractical. Credit purchases did not require an explicit contract and thus often led to a loss in sales for the retailers. However, to remain competitive, retailers allowed credit transactions so that their customers remained regular. This shows that informal credit institutions in urban local communities functioned as informal insurance networks in rural economies.
@@
This study’s primary contribution is to add analytical perspectives on consumption smoothing behavior to economic history studies. Studies of how working-class households compensated for income loss during the late 19th to the early 20th century have focused on Britain. These studies used cross-sectional variation in income to estimate the elasticities of the temporary income sources. Their detailed analyses have revealed the roles of financial institutions and labor supply adjustments in compensating for income loss in working-class households (Scott and Walker 2012). [^2] They found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use savings institutions. Horrell and Oxley (2000) highlight the important role of labor supply adjustments among family members to manage the income losses of the head in working-class households in late 19th-century Britain. Kiesling (1996) and Boyer (1997) focused on the risk-mitigation strategies among the cotton workers households in Lancashire during the recession caused by the Civil War. However, there is little evidence of how those risk-coping strategies smoothed their consumption. [^3] Other strands of literature provide supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005), building image on the historical consumers in the long-run (Francks 2009), and describe the relationship between consumption and production (Gordon 2013). Hence, this study provides a rare window into urban household financial behavior during industrialization by utilizing unique household longitudinal budget data from prewar Japan. It also implements a systematic panel data analysis to offer comparable estimates of the income elasticity for future studies.
The second contribution is revealing the roles of informal credit institutions in historical metropolis. Evidence from development economics studies indicates that the availability of informal credit institutions relates to the stability of consumption in rural economies, especially when the availability of formal financial institutions is limited. Udry (1994) showed that state-contingent loans among relatives and neighbors could partially insure against idiosyncratic shocks on rural households in northern Nigeria. A detailed field survey revealed that asymmetric information was no longer important in those private credit contracts (Udry 1990). [^4] Fafchamps and Lund (2003) also presented evidence that the households in the rural Philippines used informal networks with relatives and friends as a risk-mitigation strategy for idiosyncratic shocks. Related studies consider risk-sharing among households with similar cultural and institutional backgrounds. Grimard (1997) showed that insurance among households in the same ethnic group could partially mitigate idiosyncratic shocks in the multiethnic economy of Côte d’Ivoire. Munshi and Rosenzweig (2016) presented the importance of rural insurance networks among workers in the same caste in the context of spatial wage inequality in India. Nonetheless, there is little evidence on the function of informal credit institutions in the historical context. Ogasawara (2024) observed the role of official financial institutions, such as postal savings and pawnshops, in compensating temporary income losses in the early 20th-century Osaka. [^5] Note that Osaka was middle scale city of approximately half the size of Tokyo City circa 1920 (Online Appendix Table A.1). However, the function of informal credit institutions in consumption smoothing has not been considered due to limited data availability. By considering the roles of both formal and informal financial institutions simultaneously, this current study documents the contribution of informal credit transactions by local retailers.
Relatedly, an examination of how, and the extent to which, the local society could ensure household consumption provides valuable insights into existing urban history literature. After World War I, working-class households in Tokyo could establish and sustain their families (Nakagawa 1985). The results of this study suggest that the expansion of urbanization and credit transactions through the war underpins this traditional historical view. Another suggestion relates to the findings by Munshi and Rosenzweig (2016), who linked spatial wage inequality and men’s mobility in India with the availability of risk-mitigation networks among workers in the same caste. From this perspective, local credit institutions with lower barriers may motivate rural workers to urbanize, which may partially explain why Tokyo grew so fast circa WWI. This hypothesis may be worthy of testing in future research.
The third contribution is to offer a case study of a metropolis in a developing economy. Extant development economics studies on consumption smoothing behavior have focused particularly on the rural economies (e.g., Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002). [^6] Recent field experimental studies also focus on village economies (Comola and Prina 2023; Macours et al. 2022). While Townsend (1995), Skoufias and Quisumbing (2005), and Aguila et al. (2017) provided case studies on middle-sized contemporary economies, there is little evidence on agglomeration economies like prewar Tokyo. Urbanization is, however, now a worldwide phenomenon. The proportions of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020). This requires a deeper understanding of consumption behavior among urban working households in the economic development process. Moreover, there is little research on the role of financial institutions among the household sector in agglomeration economies. This study considers a set of homogeneous households in a representative manufacturing area in prewar Tokyo, which provides a valuable empirical setting for testing risk-sharing strategies in the historical metropolis. [^7] Focusing on a sample that shares similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). Caution is needed when comparing results, as the modern financial system differs in many ways from the past. However, my findings show that informal credit institutions in the local urban community could mitigate vulnerability among households similar to rural developing economies (Udry 1994; Fafchamps and Lund 2003).
@@ 2.2 Tsukishima: A Representative Manufacturing Area
The population in Tokyo City has been growing since the end of the 19th century, and industrialization due to WWI spurred its growth. Tokyo City’s population share increased from $2.5$% in 1889 to approximately $4$% by 1920.131313In 1920, there were more than $10,000$ municipalities in Japan, suggesting how large Tokyo was. There was a difference of about 1 million between Tokyo and Osaka, the second-largest city by population. Online Appendix Figure A.1a summarizes the trends in the number of people in the top five cities between 1889 and 1920. Throughout the 1910s, the percentage of industrial and commercial workers increased in Tokyo City, building the city as a gigantic commercial and industrial place (Tanimoto 2002, p. 8). Column 1 of Panel A in Table 1 shows that nearly half of the male workers in Tokyo worked in the industrial sector as of 1920.
@@
Panel B in Table 1 summarizes the population and occupational statistics from the census to examine the characteristics of working-class households in Tsukishima. First, the average household size in Tsukishima is $4.3$—similar to but slightly smaller than the values in Tokyo City ($4.6$) and Kyobashi Ward ($4.7$). The percentage of the married male population is $40$%, which is similar to the range of the figures for Tokyo ($38$%) and Kyobashi ($36$%). The average age of men in Tsukishima is $25.1$, slightly lower than the figures for Tokyo and Kyobashi. While the sex ratio by age group in Tsukishima shows a similar trend to that of Tokyo and Kyobashi, the ratios for the 1414141419191919 and 2020202029292929 age bins are higher than those. Overall, these statistics reflect the fact that the typical household structure is a couple with two children. The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory worker households. The percentage of married men in Tsukishima suggests that a certain proportion of the households in their early 20s might have been couples before they started raising children.
@@ 3.2 Sample Characteristics
The estimation strategy used for testing consumption-smoothing and risk-coping strategies requires panel structures for aggregate income and expenditure (Section 4). I use the semi-monthly series for the baseline analysis, as the head’s income was paid twice a month. [^30] Section 3.3 provides details on the definitions of aggregate income and expenditure. Among the $40$ THBS households, there are five cross-sectional units in our semi-monthly data. I excluded these five cross-sectional units, leaving $35$ households. Next, I removed a household with incomplete income information. Finally, I excluded a household for which no information on payments to the credit purchases was recorded. Consequently, my THBS sample for the semi-monthly series included $33$ households with panel structures from January 12 to November 11, 1919.313131I also used the adjusted monthly panel dataset for sub-analysis in Section 4. Following the same trimming steps, my adjusted monthly panel dataset included $26$ households for the same period. Note that several units have only three semi-month cells, for example, less than two adjusted months. These units have a panel structure in the semi-monthly panel dataset, but not in the adjusted monthly panel. Online Appendix B.3 summarizes the finer details of the original data structure and trimming.
@@ 3.3 Aggregation
To manage the mismatch, I aggregated the daily observations between the 12th and 26th for the first half and those between the 27th and 11th of the next month for the second half. [^40] This means that the number of days included in a semi-month ranges from 13 (i.e., in February) to 16 (e.g., January). Note that this heterogeneity in the number of days over different semi-months is a cross-sectional constant. Thus, these differences are entirely captured using the semi-month fixed effects in the regression. This range is based on the fact that paydays in the second half of a month measured in the THBS dataset ranged from 27th to 31st. [^41] Note that I do not use the definition of semi-month starting from the 2nd (to the 16th) in each month because this method of stratification places the payday at the end of each semi-month cell. Online Appendix B.7 provides the evidence that while the definition using 2nd in each month as the threshold provides an improved series in income, the expenditure cannot chase the income correctly because the expenses measured under this definition are based on the income in the previous semi-month cell. My preferred definition was using the 12th of each month as this threshold offers a clean series of income and expenditure. Figure 4c shows the time-series plots of the income and expenditure from this adjusted semi-monthly panel dataset, whereas Figure 4e illustrates the unadjusted (calender) semi-monthly series using the 15th of each month as the threshold of the semi-month. Clearly, Figure 4e provides very rough measures of income and expenditure due to the measurement errors from the maladjustment of paydays. The most significant improvement in Figure 4c is that the expenditure can chase the income in each month, reflecting the households’ consumption strategies based on their income measured in the approximate exact timing. Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which is unavailable under the unadjusted series (Figure 4f). Online Appendix B.8 shows that these measurement errors in the unadjusted dataset strongly attenuate the estimates.